The points A,B and C lie on the same straight line, the distance between AB being 60m. A particle crosses the point A with the velocity 2m/s and moves with uniform acceleration along the straight line. After crossing the point B, it takes 10 sec to reach C with velocity 5m/s. Find the acceleration of the particle.
1
Expert's answer
2019-08-12T15:41:18-0400

Let

acceleration of the particle = a

time to travel from A to B = t

speed at point A = VA

speed at point C = VC

According to constant acceleration motion formulas


"|AB| = V_A\cdot t+a\frac{t^2}{2}""a=\frac{V_C-V_A}{t+10}""60=2t+a\frac{t^2}{2}\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (1)""a=\frac{5-2}{t+10}=\frac{3}{t+10}\ \ \ (2)"

Substituting a from (2) into equation (1):


"60=2t+\frac{3t^2}{2(t+10)}""120(t+10)=4t(t+10)+3t^2""120t+1200=4t^2+40t+3t^2""7t^2-80t-1200=0"

using quadratic formula:


"t_{1,2}=\frac{-(-40)\pm \sqrt{40^2-7\cdot(-1200)}}{7}=""\frac{40\pm \sqrt{10000}}{7}=\frac{40\pm 100}{7}"

"t_1,=20,\ t_2=-60/7"

Since t > 0 t = t1 = 20s

From equation (2)

"a=\frac{3}{20+10}=\frac{3}{30}=\frac{1}{10}m/s^2"

Answer:

"\frac{1}{10}m/s^2"